OpenStudy (anonymous):

How do you solve? x+"square root" of x +6

OpenStudy (anonymous):

\[x+\sqrt{x}+6\]

OpenStudy (anonymous):

\[x+6=\sqrt x\] squaring both the sides \[x^2=(x+6)^2\]\[x^2=x^2+12x+36, 12x+36=0\]now you can do it ;)

OpenStudy (anonymous):

thank you for the help